OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -2 and a5 = -512?

OpenStudy (anonymous):

\[a5= a1* r^4\\ -512= -2 r^4\\ r=4 \\or \\r=-4 \]

OpenStudy (anonymous):

if r=4 \[ a_7=-2 (4)^6=-8192 \]

OpenStudy (anonymous):

We obtain the same a7 for r=-4